OpenStudy (anonymous):

In one experiment, 0.847 moles of NO is mixed with .492 of O2. Determine which of the two reactants is the limiting reactant. Calculate also the number of moles of NO2 produced. 2NO + O2 -> 2NO2

OpenStudy (aaronq):

divide the number of moles of each species by it's stoichiometric coefficient to find the amount of "normalized" moles. the one less "normalized" moles is the limiting reactant. for the rest of the problem use the number of moles of the limiting reactant. Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

I don't understand. what is nA and nB and nC?

OpenStudy (aaronq):

n = moles, \(n_A\) = moles of species A